### v1

# Towards Reducible Uncertainty Modeling for Reliable Large Language Model Agents

Abstract
Uncertainty quantification (UQ) for large language models (LLMs) is a key building block for safety guardrails of daily LLM applications. Yet, even as LLM agents are increasingly deployed in highly complex tasks, most UQ research still centers on single-turn question-answering. We argue that UQ research must shift to realistic settings with interactive agents, and that a new principled framework for agent UQ is needed. This paper presents the first general formulation of agent UQ that subsumes broad classes of existing UQ setups. Under this formulation, we show that prior works implicitly treat LLM UQ as an uncertainty accumulation process, a viewpoint that breaks down for interactive agents in an open world. In contrast, we propose a novel perspective, a conditional uncertainty reduction process, that explicitly models reducible uncertainty over an agent’s trajectory by highlighting “interactivity” of actions. From this perspective, we outline a conceptual framework to provide actionable guidance for designing UQ in LLM agent setups. Finally, we conclude with practical implications of the agent UQ in frontier LLM development and domain-specific applications, as well as open remaining problems.
[…]
1 Introduction
Figure 1: Comparison between UQ setups. Traditional LLM UQ (a) measures the uncertainty of an answer given a question, whereas the UQ for LLM reasoning (b) expands the problem by considering multi-step responses rather than a single response. Agent UQ (c) goes further by considering continual interactions between agent and user/environment across the trajectory, making it a multi-turn, interactive inference setup (example sourced from $\tau$-bench Airline; barres2025tau).
LLM-based agents operating in open-world environments take actions that have real consequences: making costly bookings, modifying databases, or issuing irreversible commands [openai2025agent, google2025agent] at the next level of autonomy [Steinberger2026OpenClaw]. In such settings, failures are no longer limited to incorrect text generation: agents may act prematurely under unresolved ambiguity, propagate errors across long interaction trajectories, or commit to outcomes that are costly or difficult to undo. To be deployed responsibly, agents must be able to assess and act upon the likelihood of failure [kochenderfer2015decision]. This makes reliable uncertainty quantification (UQ) an urgent and central requirement for agentic systems.
Despite the importance, most existing UQ research treats LLMs as static oracles: a system examined in isolation, prompted once, and evaluated by the uncertainty of a single response or a chain of responses (Figure 1, left). These methods implicitly assume a static system in which no new information is acquired after the initial prompt [malinin2021uncertainty, kadavath2022language, farquhar2024detecting]. Under this assumption, uncertainty is treated as a point-wise estimate or uni-directional propagation. While appropriate for single-turn inference and non-interactive multi-step reasoning, they do not meet the growing needs of UQ in agentic AI settings featured by its long-horizon interactive environment.
Interactive agents operate in multi-turn interactive environments with feedback from users, tools, and external systems (Figure 1, right). For example, a flight-booking agent may need to decide whether to finalize a reservation or ask follow-up questions about dates, budget, or layover preferences. Early uncertainty should prompt information-seeking actions, while subsequent dialogue and database queries can resolve ambiguities and reduce uncertainty, enabling confident execution. Here, uncertainty does not strictly accumulate–it may also decrease as a direct consequence of interaction, even as the trajectory grows longer.
In this paper, we argue that making LLM agents trustworthy requires a shift in focus: from pointwise uncertainty of answers to structured uncertainty flow in open, interactive decision processes; and a new framework is necessary for this. We present the following perspective:
Uncertainty quantification for an LLM agent must incorporate interactivity to model reducible uncertainty: in non-isolated, agentic systems, total uncertainty should be allowed to decrease as new information is acquired through interaction.
To ground this perspective in a formal discussion, we first provide a concrete definition of agent UQ and a general formulation that captures broad categories of existing UQ setups (Sec. 3). Specifically, we model the agent’s problem-solving trajectory as a stochastic process over actions, observations, and environment states, represented as a simple graphical model. Then, we define both turn-level and trajectory-level uncertainty in a simple, probabilistic expression, where we show that many prior UQ approaches arise as special cases.
This unified view also exposes a fundamental limitation of existing methods: uncertainty is implicitly treated as accumulating monotonically along the trajectory, regardless of the role played by different actions; thus fails to relate uncertainty to a task reward. To address this limitation, we introduce in Sec. 4 a conceptual framework that models agent UQ as a conditional uncertainty reduction process, through an information-gating mechanism that operationalizes the distinction between uncertainty-increasing and uncertainty-reducing actions by considering their interactivity. We further derive the analytic bounds that improve the interpretability of the resulting uncertainty estimates.
In addition, Sec. 5 discusses practical implications of agent UQ across diverse domains, including healthcare, software engineering, and robotics, as well as frontier LLM research on adaptive reasoning and post-training. Finally, we outline open problems unique to uncertainty modeling in agentic systems (Sec. 6). Together, we view this work as a step toward providing the field with a necessary foundation and clearer direction for future research on uncertainty-aware agentic systems.
  • We formalize agent UQ as a distinct problem setup and show why existing LLM UQ formulations are insufficient for interactive agent setups.
  • To tackle that limitation, we present a conceptual framework for agent UQ by viewing it as a conditional uncertainty reduction process.
  • To give actionable insights and promote discussion, we describe promising examples of agent UQ applications in various domains and also remaining open problems.
2 Related Work
UQ in LLMs.
In contrast to classic UQ [neal1992bayesian, mackay1992practical, gal2016dropout], LLM UQ brings extra challenges from its computational cost and free-form outputs. Common approaches include aggregating output token probability [zhang2025token, marina2020unsupervised, malinin2021uncertainty, zhang2023enhancing, fadeeva2024fact, duan2024shifting, aichberger2024rethinking], measuring consistency over samples [manakul2023selfcheckgpt, farquhar2024detecting, nikitin2024kernel], verbalizing confidence [kadavath2022language, lin2022teaching, yona2024can], or conformal prediction [kumar2023conformal, cherian2024large, quach2024conformal]. While promising, they mostly use single-step question-answering as a testbed, which limits their practical applicability to agentic setups.
Figure 2: Graphical model for an agent problem-solving trajectory with examples. Given a task specification $E_{0}$ and an initial user query $O_{0}$, an agent spans a multi-turn trajectory characterized by a chain of action AA, observation $O$, and environment state $E$. This simple abstraction describes some representative agentic prompting methods such as ReAct [yao2022react]. See Appendix A.1 for details.
UQ in reasoning and agent setups.
Beyond single-step generation, researchers show growing interest in measuring the confidence of LLMs during their reasoning. This kind of (un)certainty information has been used to improve the quality of response [wang2023selfconsistency, kang2025scalable, fu2025deep] or to guide a model during inference or training [hu2024uncertainty, lightman2024lets, zhu2025uncertainty, wang2025beyond, wang2025harnessing, cheng2025reasoning]. Recently, there have also been attempts to model uncertainty in interactive and/or agentic setups [han2024towards, zhao2025uncertainty, duan2025uprop, frankel2024conformal, chan2025conformal]. Although UQ here aims to model uncertainty dynamics over long-horizon generation, existing methods do not capture the reducible nature of uncertainty in open environments [kirchhof2025position], highlighting the need for a new framework for agent UQ.
3 A General Formulation for Agent UQ
3.1 Problem Statement and Definitions
Imagine a flight booking agent depicted in Figure 1 right. Here, the types of action $A$ and observation $O$ are diverse: for $A$, there is information-seeking, questioning for clarification, calling an API function, and thinking or updating without interactive behaviors. The observation $O$, on the other hand, can be a database status report, tool-calling results, or the user’s clarification. Both action and observation can be displayed in flexible forms of natural language messages or structured strings, e.g., code.
Concretely, given a task specification $E_{0}$, e.g., a flight booking, a user makes a stochastic initial query $O_{0}$, and the agent first requests the user’s ID to get relevant information. Then, it calls a programmed function to get privileged information from the database. Not only does the agent interact with the user or tool, but it also thinks about better future actions (i.e., planning) or makes a system update without external observation $O=\{\varnothing\}$. All interactions and other actions are stored in environment state $E$ for context-grounded actions over a long horizon inference. This makes up a trajectory $\mathcal{F}_{\leq t}$ (chain of $A$, $E$, and $O$) up to $t$-turn which continues until the task-specific termination rule is met, e.g., when the agent submits an answer [yang2018hotpotqa, jimenez2024swebench, wei2025browsecomp], the user generates a stop sign [yao2024tau, barres2025tau], or goal completion [shridhar2021alfworld] with multiple constraints [xie2024travelplanner].
Now, to formalize UQ in this real-world agent context, we begin with a concrete definition of the agent system in Definition 1, and introduce a graphical model for the agent trajectory in Figure 2. This can generalize many existing agent benchmark setups [mialon2023gaia, yao2024tau, yang2018hotpotqa, song2023llm, jimenez2024swebench, xie2024travelplanner, liu2024agentbench, wei2025browsecomp].
Definition 1 (Stochastic Agent System).
Let $E_{i}$ be an environment state, a mixture of a context memory of agent-user-tool interaction logs up to the $i$-th turn that is fully accessible to the agent, and a state of the system database, which is not directly accessible to the agent (only partially observable via tool-calling). Let $O_{i}$ and $A_{i}$ be the $i$-th turn observation and action derived from distributions $P$ and $P_{\pi,\mathcal{T}}$, respectively, where $\mathcal{T}$ indicates a tool set and $\pi$ is an LLM policy. Then, under a task specification $E_{0}$ and the user’s initial query $O_{0}$, we define the trajectory as a sequence of turns $\mathcal{F}_{\leq T}=\{(A_{t},E_{t},O_{t})\}_{t=0}^{T}$ where $\mathcal{F}_{0}=(E_{0},O_{0})$. Given a deterministic update function $h(\cdot)$, the generative process for each step $t$ is defined as: $A_{i}$ $\sim P_{\pi,\mathcal{T}}(\cdot|E_{i-1},O_{i-1}),\penalty 10000\ \penalty 10000\ O_{i}\sim P(\cdot|A_{i},E_{i}),\penalty 10000\ \penalty 10000\ E_{i}=h(E_{i-1},O_{i-1},A_{i}).$
To express this trajectory with a simple mathematical representation, we introduce a dynamic Bayesian network [murphy2002dynamic] in Figure 2. With this graphical model, dependency structures of $A$, $E$, and $O$ can be explained as follows: (1) since the environment state $E_{i-1}$ contains the entire conversation history of $A_{\leq i-1}$ and $O_{<i-1}$ up to the i1i-1 step [^1] An adaptive memory structure [xu2025mem, xiong2025memory] can replace this ever-growing memory., the current action $A_{i}$ depends only on $E_{i-1}$ and $O_{i-1}$; (2) the current observation $O_{i}$ from the environment also depends only on $A_{i}$ and $E_{i}$; (3) the transition of the environment state ($E_{i-1}\rightarrow E_{i}$) can be defined as a deterministic function [^2] In general, this system transition can be also stochastic, but we consider a simplified yet still realistic scenario. $h(\cdot)$ which stacks the conversation logs up in a memory module and conducts a rule-based modification (write, update, or delete) on a previous system state given the current action.
Under this Bayesian network, the joint probability for the trajectory can be factorized as below,
$P(\mathcal{F}_{\leq T})$ $=P(E_{0},O_{0})\Pi_{i=1}^{T}P(\mathcal{F}_{i}|\mathcal{F}_{i-1})$
$=P(E_{0},O_{0})\Pi_{i=1}^{T}P_{\pi,\mathcal{T}}(A_{i}|E_{i-1},O_{i-1})P(O_{i}|A_{i},E_{i}),$
where $\mathcal{F}_{0}=(E_{0},O_{0})$, and $P(E_{i}|\cdot)=1$ as it is determistic. Now, we define the UQ problem for this stochastic agent system as follows.
Definition 2 (Agent UQ).
Let $U(X)\geq 0$ be an uncertainty function that takes a random variable $X$ or its distribution to produce a non-negative real value. Given a $T$-turn trajectory, agent UQ aims to estimate both a turn-level uncertainty $U(\mathcal{F}_{t}|\mathcal{F}_{t-1})$ for t=1,,Tt=1,...,T and a trajectory-level uncertainty $U(\mathcal{F}_{\leq T})=U(\mathcal{F}_{0},...,\mathcal{F}_{T})$, as a joint total uncertainty.
In the machine learning field, the most popular choices of the uncertainty function $U(X)$ are Shannon entropy [shannon1948mathematical] $\mathbb{E}[-\log P(X)]$, negative log probability $\log P(X=x)$, and their variants. These information-theoretic measures of uncertainty have a nice property, chain-rule [cover1999elements], making them suit for sequential modeling. For example, given $U(\mathcal{F}_{t}):=\mathbb{E}[-\log P(\mathcal{F}_{t})]$, we have the following uncertainty expansion, allowing us to model the agent’s total uncertainty as a simple arithmetic of component-level uncertainties  [^3] See Appendix A and B for a derivation and alternative measures, e.g., informational energy [pardo1991information] and nonextensive entropy [gell2004nonextensive].:
$U(\mathcal{F}_{\leq T})$ $=U(E_{0},O_{0})+\sum_{i=1}^{T}U(\mathcal{F}_{i}|\mathcal{F}_{i-1})$
$=U(E_{0},O_{0})+\sum_{i=1}^{T}[U(A_{i}|E_{i-1},O_{i-1})+U(O_{i}|A_{i},E_{i})],$
where $U(E_{0},O_{0})=U(O_{0}|E_{0})+U(E_{0})$ denotes an initial query uncertainty plus a task volatility.
Desideratum.
While previous LLM UQ setups usually focus on conditional uncertainty of a final answer, agent UQ aims at estimating joint uncertainty over the full trajectory, where the estimate from a calibrated agent should be predictive of the reward on that trajectory. Formally, for all $r_{1}>r_{2}$,
$\mathbb{E}\big[U(\mathcal{F}_{\leq T})|r(\mathcal{F}_{\leq T})=r_{1}\big]<\mathbb{E}\big[U(\mathcal{F}_{\leq T})|r(\mathcal{F}_{\leq T})=r_{2}\big],$ (1)
where $r(\cdot)$ denotes a real-value reward function commonly set as a success-failure binary annotator.
We draw analogies between agent UQ and probabilistic Turing machines [gill1974computational], as well as belief tracking in Partially Observable Markov Decision Processes [kaelbling1998planning], highlighting that agent UQ is grounded while still having its unique edge in contrast to these existing theoretical models in terms of setups and focus. See Appendix C for discussion.
3.2 A Unified View on Existing UQ Setups
We now show that our agent UQ formulation captures various existing UQs as special cases.
Reduction to single-step LLM UQ.
When $t=1$ and the action space is restricted to responses, LLM UQ can be cast as a special case
$U(\mathcal{F}_{\leq T})=U(O_{0})+U(A_{1}|O_{0})\geq U(A_{1}|O_{0}),$
where $O_{0}$ and $A_{1}$ denote the given question and the corresponding model answer, respectively. Most current LLM UQ literature addresses only the conditional answer uncertainty $U(A_{1}|O_{0})$ [malinin2021uncertainty, fadeeva2023lm, farquhar2024detecting]. A few also consider the question ambiguity $U(O_{0})$ [hou24decomposing, tomov2025the] to reflect the reality of wild user query [min2020ambigqa].
Reduction to multi-step reasoning UQ.
When the action space involves multi-step reasoning (e.g., chain-of-thought, wei2022chain), we have
$U(\mathcal{F}_{\leq T})$ $=U(O_{0})+\sum_{i=1}^{T}U(A_{i}|A_{<i},O_{0})$ (2)
$\geq\sum_{i=1}^{T}U(A_{i}|A_{<i},O_{0})$ (3)
$\geq\max_{i\in\{1,...,T\}}U(A_{i}|A_{<i},O_{0})$ (4)
$\geq\sum_{i=1}^{T}w_{i}U(A_{i}|A_{<i},O_{0}),$ (5)
where $w_{i}\geq 0$ and $\sum_{i=1}^{T}w_{i}=1$. Here, $O_{0}$ and $A_{T}$ denote the given initial query and the model’s final response, and $A_{i},...,A_{T-1}$ are intermediate outputs. Similar to the single-step setup, most existing multi-step UQs [tanneru2024quantifying, tao2025revisiting] do not consider the question ambiguity $U(O_{0})$, so that reduces Eq. 2 to Eq. 3, while a few consider it [leang2025picsar]. Meanwhile, rather than summing up all the step-wise uncertainties, some approaches consider the most uncertain one, i.e., the lowest confidence used in fu2025deep, which can be represented by Eq. 4. Finally, Eq. 5 expresses various weighted average methods, such as a standard length-normalized uncertainty [kang2025scalable] by setting $w_{i}=\frac{1}{T}$, and a tail confidence [fu2025deep] by setting $w_{T}=1$, and other clever weighting strategies, i.e., top-K mean [fu2025deep] or model-based estimates [zhao2025uncertainty].
Connection to process reward modeling.
We further draw a connection between this trajectory-level UQ problem and the process reward modeling problem [uesato2022solving, lightman2024lets, li2025pqm] where we assign credit (reward) for each intermediate reasoning step, and then it can be utilized to evaluate the reasoning trajectory by aggregating the rewards across steps to derive a trajectory-level reward. For example, lightman2024lets proposed the product of the step-wise probabilities and the minimum probability between steps as the aggregation method. These are analogs to Eq. 3 and Eq. 4, respectively.
Figure 3: Limitation of existing UQ (a) and our suggestion (b). Prior works just concern the evidentiality when designing or evaluating UQ while neglecting interactivity, so they may fail to reliably capture the agent’s failure. We urge expanding a dimension of interest, interactivity, and moving on to reducible uncertainty modeling of agents.
Limitation of existing frameworks.
From this unified view, we see that they all model the multi-step uncertainty as a uni-directional propagation, i.e., every step-level uncertainty positively contributes to the total uncertainty, without distinguishing action types at each turn. One might claim that the average uncertainty $\frac{1}{T}\sum^{T}_{t=1}U(\mathcal{F}_{t}|\mathcal{F}_{t-1})$ can already be fine, as it might produce a lower value for the chain of certain actions and vice versa. However, it does not consider the type of action which is critical in agentic setups, where an agent’s interactivity is key to earning a reward. See Fig. 3 (a), a scenario wherein a user does not fully specify their goal upfront. After searching for options, the agent either reasons itself or interacts with the user to delegate decisions. Although both trajectories are evidential, interaction-oriented ones tend to achieve higher rewards, despite reasoning-oriented ones often yielding greater confidence. Yet, existing methods can produce misleading uncertainty estimates, fail to meet the desiderata (Eq 1).
4 How to Model Uncertainty Dynamics of Interactive Agents?
4.1 Unique Challenge
In contrast to the LLM UQ, agent UQ spans a multi-turn trajectory in an open environment. This presents a unique challenge: modeling cascade uncertainty in an interactive system. Here, the desired UQ method should be able to precisely aggregate uncertainties across the entire trajectory and account for additional information gained through interactions. While there is emerging research on step-wise uncertainty aggregation [kang2025scalable, fu2025deep], they examine an isolated, noninteractive setting where an agent does not receive any external feedback during inference (e.g., Fig. 1, b). Meanwhile, agent UQ considers a setup full of interactions with a user and an environment, making it a non-isolated, interactive system.
Since an agent in this open world can gather additional information, make clarifications, and request confirmations at each step by interacting with the user and external databases through conversations and tool calls, the uncertainty dynamics in this interactive system includes not only propagations of turn-level uncertainty, but also reductions. As shown in Fig. 3, the existing perspective does not consider “interactivity” in agent UQ and prohibits reducible uncertainty modeling. In this paper, we argue that we must introduce interactivity in agent UQ to model bi-directional uncertainty flow.
4.2 Agent UQ as a Conditional Uncertainty Reduction Process
We need a new perspective to model an agent’s uncertainty dynamics that accounts for both reductions and increases, depending on the agent’s interactivity at each step. While there are multiple ways to realize this concept, we propose one possible implementation as a prototype: information gating, which goes hand in hand with the information-theoretic measure of uncertainty noted in Sec. 3.
Specifically, we derive a lower bound of total uncertainty $U(\mathcal{F}_{\leq T})$ that admits a selective reduction or increase in uncertainty throughout the trajectory:
$U(\mathcal{F}_{\leq{T}})=U(E_{0},O_{0})+\sum_{i=1}^{T}U(\mathcal{F}_{i}|\mathcal{F}_{i-1})$ $=U(E_{0},O_{0})+\sum_{i=1}^{T}[U(A_{i}|E_{i-1},O_{i-1})+U(O_{i}|A_{i},E_{i})]$
$=U(E_{0},O_{0})+\sum_{i=1}^{T}U(A_{i}|E_{i-1},O_{i-1})[1+\frac{U(O_{i}|A_{i},E_{i})}{U(A_{i}|E_{i-1},O_{i-1})}]$
$\geq U(E_{0},O_{0})+\sum_{i=1}^{T}U(A_{i}|E_{i-1},O_{i-1})g(\mathcal{F}_{i}).$ (6)
Here, we introduce a conditional gating function $g(\cdot)$ that reduces or increases a current turn uncertainty depending on the action. That is, for a set of valid uncertainty reduction actions, $\mathcal{A}^{-}$, we define:
$g(\mathcal{F}_{i})=\begin{cases}\frac{-\text{Info}(O_{i};O_{0}|E_{i}\backslash O_{0})}{U(A_{i}|E_{i-1},O_{i-1})}&\text{if}\penalty 10000\ A_{i}\in\mathcal{A}^{-},\\ 1+\frac{U(O_{i}|A_{i},E_{i})}{U(A_{i}|E_{i-1},O_{i-1})}&\text{otherwise.}\end{cases}$
where $\text{Info}(\cdot;\cdot)$ denotes mutual information or pointwise mutual information that measures the amount of information gain by having the current observation $O_{i}$ w.r.t. initial query $O_{0}$ given conversation history $E_{i}\backslash O_{0}$. For each turn, if the agent’s action is classified as an interactive (inviting a user or a tool) and evidential one (factual or not in conflict with stored data), then $g(\cdot)$ produces the amount of information gain with negative sign to realize uncertainty reduction; otherwise, it propagates uncertainty (See Fig. 4).
Figure 4: Illustration on the proposed agent UQ paradigm. We propose the conditional uncertainty reduction process for LLM agents by discerning interactive and evidential actions from others.
Importantly, the signed gating function handles uncertainty in a more interpretable manner than traditional estimates by letting not only quantitative but also directional interpretations. Besides, this uncertainty model allows us to derive the analytic extrema, i.e., maximum and minimum, of total uncertainty in closed forms as shown in Lemma 1.
Action Category Interactivity Example
Information-gathering interactive Agent uses a read tool (e.g., get_reservation_details or search_flights) to retrieve data or asks the user for missing information, e.g., “Could you provide your reservation number and last name?”.
Asking clarification or confirmation to user interactive Agent asks the user preferences or confirmation on decision, such as, “Do you want me to proceed with booking FQ8APE?”.
Thinking non-interactive Agent plans the future action sequence based on the previous trajectory.
State-changing tool call (writing) non-interactive Agent calls the tools that modify the database (e.g., cancel_reservation, book_reservation, update_reservation_flights) commit to an outcome.
Providing final information to user non-interactive Agent reports the result of an action, such as “Your reservation has been cancelled; your refund will be processed” or “Your flight has been rebooked to SFO departing at 8 a.m.”.
Table 1: Example action classes in Airline booking scenario in $\tau$-bench. A number of the agent’s actions can be categorized into five classes, which may reduce or increase the total uncertainty. See Appendix E for extended analyses.
Lemma 1 (Extrema of Information Gating).
Let the lower bound of agent total uncertainty in Eq. 6 be $\tilde{U}(\mathcal{F}_{\leq T})$, denote $U(X):=H(X)=\mathbb{E}[-\log P(X)]$ and $\text{Info}(X;Y):=I(X;Y)=\mathbb{E}[\log\frac{P(X,Y)}{P(X)P(Y)}]$, then, we have:
$\tilde{U}(\mathcal{F}_{\leq T})$ $\geq H(E_{0},O_{0})-\sum_{i=1}^{T-1}I(O_{i},O_{0}|E_{i}\backslash O_{0})+H(A_{T}|E_{T-1},O_{T-1}),$
$\tilde{U}(\mathcal{F}_{\leq T})$ $\leq H(E_{0},O_{0})+\sum_{i=1}^{T}H(A_{i},O_{i}|E_{i-1},O_{i-1}).$
Proof.
If $A_{i}\in\mathcal{A}^{-}$ (resp. $A_{i}\notin\mathcal{A}^{-}$) for all $1\leq i\leq T-1$, then $\tilde{U}(\mathcal{F}_{\leq T})$ becomes a monotonic uncertainty reduction (resp. propagation) process, deriving above inequalities (See Appendix F for details). ∎
Interpretation.
Lemma 1 states that an agent’s total uncertainty can be expressed as an ever-decreasing process starting from a task-specific initial uncertainty reduced by the amount of extra information delivered by current observation, given the previous trajectory till its final behavior in the ideal case; whereas it is continually stacked up by the amount of joint entropy between action and observation per turn in worst case. Together with the directional interpretation turn-level uncertainty (increase or decrease), these analytic extrema help us better understand the numerical uncertainty, which is usually hard to interpret [van2019communicating, belem2024perceptions, kirchhof2025self].
Sketch of implementation.
To operationalize this idea, we first need an action classifier to determine whether $A_{i}\in\mathcal{A}^{-}$ or not. One can adopt a model-based evaluator [kim2025beyond] and a rule-based verifier [zeng2025reinforcing] to assess the interactivity and evidentiality of actions, respectively, where $A_{i}\in\mathcal{A}^{-}$ if and only if $A_{i}$ is both evidential and interactive. An example categorization of actions in an airline assistant task is provided in Table 1. After that, four different quantities need to be estimated: (1) initial query uncertainty, (2) action uncertainty, (3) observation uncertainty, and (4) mutual information between the query and observation given the previous trajectory. For these estimations, we can leverage the output probability [malinin2021uncertainty, aichberger2024rethinking], consistency in multiple trials [farquhar2024detecting], world modeling [chae2025web, anonymous2026dreamphase], and a learning neural estimator [molavipour2020conditional, gritsai2025mist]. We leave the actual implementation for future work. See Appendix D for additional notes.
5 Practical Implications
Developing an agent UQ framework is not merely a theoretical exercise but a prerequisite for deploying LLM agents in non-deterministic real environments. We outline implications for frontier LLM research and three specialized domains: healthcare, coding, and robotics, in which agent UQ may have profound downstream effects, thereby incentivizing policymakers, practitioners, and researchers.
5.1 Advancing Frontier LLMs
Test-time scaling through reasoning is now a de facto standard technique to enhance the quality of LLM responses on challenging tasks [wei2022chain, muennighoff2025s1]. However, there is still huge room for improvement, e.g., mitigating overthinking and/or pursuing efficiency [sui2025stop, aggarwal2025optimalthinkingbench]. Agent UQ provides concrete guidance to develop an adaptive reasoning method, i.e., early stopping or interaction invoking, via uncertainty budgeting. In addition, multi-turn reinforcement learning (RL) has emerged as a promising yet challenging research theme in eliciting the agentic behavior and reasoning of LLMs [zhou2024archer, shani2024multi, zeng2025reinforcing]. Agent UQ helps tackle the main bottlenecks, credit assignment and when-to-explore problems [thrun1992efficient, pecka2014safe, choudhury2025process], limited positive feedback [lee2025banel, lee2026semi], and inspires an information pursuit policy [chattopadhyay2022interpretable] by leveraging directional and quantitative uncertainty information as feedback.
5.2 Reliable Clinical Decision Supporting Agents
Modern LLM agents offer a promising vision for healthcare, moving beyond static knowledge retrieval to autonomous systems with complex reasoning and tool-execution capabilities [wang2025survey]. However, the transition from promising prototypes to reliable clinical agents faces critical obstacles stemming from the system’s inability to recognize and manage its own limitations. We believe that agent UQ can fill this gap. For example, although ferber2025development observed a huge potential of tool-calling LLM agents in oncology diagnoses, the remaining error rate underscores the necessity of a “human-in-the-loop” workflow. Here, agent UQ can serve as a gatekeeper, e.g., flagging highly accumulated uncertainty moments to invite humans, automating during low-uncertainty periods, while pursuing total uncertainty reduction for the right final diagnosis. This aligns with the concept: adaptive healthcare that envisions a paradigm shift from passive prediction to active, risk-aware inquiry [hinostroza2025ai].
5.3 Reliable Software Engineering Agents
Although advanced coding agents [anthropic2025claude45systemcard, openai_gpt52codex2025] have pushed performance on standard benchmarks [jimenez2024swebench] to an impressive degree, the disparity between passing a controlled test and safely modifying a production-level codebase stresses the urgent need for an agent UQ framework. Agent UQ enables a more controllable exploration and commitment mechanism, wherein the agentic system knows uncertainty over bug patch candidates and decides whether to gather more evidence, e.g., check more files, or commit to a fix. Besides, considering uncertainty-triggered rollback and branching algorithm, in which an agent treats increasing uncertainty after an action as a signal to revert or branch by mirroring the checkpoint-rollback workflows of human engineers. These highlight the potential of UQ for reliable agentic coders.
5.4 Relible Embodied Agents in Cyber-Physical Systems
Embodied agents in cyber-physical systems face heterogeneous uncertainty from sensing, dynamics, and human intent [li2024embodied, fung2025embodied, team2025gemini], and they actively reduce uncertainty through interaction by re-sensing the environment or interacting with humans before committing to irreversible physical actions [xu2025compliant]. For example, a robot instructed to retrieve a fragile object may explore or seek clarification rather than execute a grasp under high uncertainty. This makes embodied autonomy a canonical instance of a conditional uncertainty reduction process, where information-gathering actions reduce uncertainty and state-changing actions increase commitment and risk. By explicitly quantifying and managing this uncertainty dynamic, our agent UQ solution offers safer action selection [zhang2025safevla], principled deferral [hagenow2025realm], and robust execution [romer2025failure].
6 Open Problems
Although we have so far presented a conceptual roadmap and the practical implications of agent UQ, significant open problems remain as follows.
Developing rigid benchmarks at scale.
As in the case of process reward modeling mentioned in Section 3.2, annotating across long-horizon trajectories introduces non-trivial efforts and costs [choudhury2025process]. Moreover, as true uncertainty is unmeasurable, i.e., only a proxy-based evaluation is possible, establishing benchmarks for agent UQ stands on the next level of difficulty than the static LLM evaluation [ielanskyi2025addressing]. Investigation on defining proper agent performance measures [kapoor2025holistic, hendrycks2025definition] with scalable evaluation and annotation methods for agent rollouts [yang2024react, chang2024agentboard, kim2025beyond] would be a crucial stepping stone.
Long-horizon black-box uncertainty estimation.
LLM agents are mostly offered by proprietary services, where we cannot access the output probability of the agent. While model owners and providers can leverage low-cost white-box UQ approaches, others may have to rely on expensive black-box UQ that requires heavy sampling per action over a long trajectory [zhang2024luq]. This poses a great challenge for low-resource users in developing UQ with state-of-the-art agents.
Solution multiplicity.
Is high uncertainty per step due to the agent’s lack of knowledge about the right action for that step? Or, is it due to the intrinsic multiplicity of valid actions in that step? We cannot identify which of them is the right source of high uncertainty! Besides, in contrast to the classic question answering, the task bid to an agent is too complex to be perfectly specified at once. Thus, each intermediate step is intrinsically ambiguous and allows multiple valid actions, making it impossible to specify the source of the high uncertainty.
UQ in multi-agent or self-evolving agent systems.
While we confine our setup to single-agent, non-evolving agent systems, there has been remarkable advancement in multi-agent systems [guo2024large, wu2024autogen] and self-evolving agents [wang2024voyager, jiang2025adaptation]. Modeling the joint uncertainty dynamics of multiple agents or considering the non-stationary toolkit and agent parameters over time brings extra challenges that may require deliberation and other formal toolkits [tsallis2004nonextensive, j2016generalising, cheng2021general].
7 Conclusion
We argue that LLM UQ research should move towards a more realistic setup, i.e., agentic inference, featuring the interactive long-horizon task episode, which requires a paradigm shift from a point-wise estimate to the bi-directional dynamics model of uncertainty. We laid the foundation for this by providing a concrete definition and general formulation of agent UQ that abstracts a wide variety of existing UQ approaches. To tackle the limitations of existing framing, we presented a novel perspective, conditional uncertainty reduction process, with actionable guidelines and a theoretical analysis. As a concluding remark, we quote Schrödinger’s note on living in physics [schrodinger1974what]: “Every living means an increase in entropy of the part of the world, and it can only survive by continually drawing negative entropy from its environment.”
Limitations
Although this paper provided a novel high-level vision for agent UQ with a crystallized formulation, we built this vision on a relatively simple setup with some assumptions: reliable observation $O_{i}$ from the environment and user, and deterministic environment update $E_{i}\rightarrow E_{i+1}$. However, agentic systems in the wild may encounter unreliable (even adversarial; kang2025trap) observations and a stochastically evolving environment, where our current proposal does not ensure its effectiveness. In that scenario, we may need to explicitly model uncertainty in the agent’s evidence stream as well, e.g., trust/ignorance about tool outputs, retrieval, and user-provided facts, by leveraging formalisms, such as subjective logic [jsang2018subjective, cheng2020there], imprecise probability theory [walley1991statistical], and formal belief representation [formalbelief22], to safely manage memory and act under unreliable observations in stochastic world.
Ethical Considerations
This work aims to establish a foundation for uncertainty quantification of LLM agents, which can directly contribute to society, science, and human well-being [han2023recipe, ren2025towards, oh2025flex, hendrycks2025definition]. We believe that the realization of our proposed framework can significantly boost the reliability of the LLM agent in a dynamic, interactive inference setup, thereby facilitating risk mitigation in high-stakes decision-making. However, we also acknowledge that the proposed framework can be misused by allowing adversaries to efficiently steer the agent towards its blind spots or manipulate uncertainty thresholds to bypass safety filters [wang2025agentvigil]. Therefore, we recommend alleviations such as limiting exposure to attack-enabling signals, red-teaming, and monitoring for abuse, which avoid overpromising what low uncertainty implies.
Acknowledgments
We sincerely thank Artem Shelmanov, Shawn Im, Hyeong Kyu Choi, Jongwon Jeong, Eunsu Kim, Mingyu Kim, Ayoung Lee, JungEun Kim, Hayun Lee, and Kyungwoo Song for their sharp feedback on the draft. This work is supported in part by the AFOSR Young Investigator Program under award number FA9550-23-1-0184, National Science Foundation under awards IIS-2237037 and IIS-2331669, Office of Naval Research under grant number N00014-23- 1-2643, Schmidt Sciences Foundation, Open Philanthropy, Alfred P. Sloan Fellowship, and gifts from Google and Amazon. Paul Bogdan acknowledges the support by the National Science Foundation (NSF) under the NSF Award 2243104 under the Center for Complex Particle Systems (COMPASS), NSF Mid-Career Advancement Award BCS-2527046, U.S. Army Research Office (ARO) under Grant No. W911NF-23-1-0111, Defense Advanced Research Projects Agency (DARPA) Young Faculty Award and DARPA Director Fellowship Award, Okawa foundation award, National Institute of Health (NIH) under R01 AG 079957, and Intel faculty awards.
Appendix
We cover the following content in this Appendix.
  • (A) Additional Details on Formulation
  • (B) Alternative Uncertainty Measures
  • (C) Extended Context
  • (D) Note on Implementation
  • (E) Action Categorization
  • (F) Missing Proof
  • (G) Discussion and Future Work
Appendix A Additional Details on Formulation
A.1 Expression of various agentic prompting methods under our formulation.
In Section 3.1, we have introduced a definition of stochastic agent system (Def. 1) with a probabilistic graphical model of specific conditional dependency assumptions 2. Under this model, we can abstract some representative agentic prompting methods [ahn2022can, yao2022react, shinn2023reflexion, patil2025berkeley] with a unified language yet distinctive conditions as below.
  1. 1.
    Vanilla function calling [patil2025berkeley]: Same as Definition 1. Stochastic process of $(A_{t},E_{t},O_{t})$ without any extra conditions.
  2. 2.
    Few-shot prompting [ahn2022can]: Just augmenting initial task-specification variable as $E^{{}^{\prime}}_{0}=E_{0}\cup\{\mathcal{F}^{(i)}\}_{i=1}^{N}$ with some example rollouts $\mathcal{F}^{(i)}=\{(A_{t},E_{t},O_{t})\}_{t=1}^{T^{i}}$, which may implicitly affect the action transition probability, $P_{\pi,\mathcal{T}}(\cdot)$, by conditioning.
  3. 3.
    ReAct [yao2022react, wang2024executable]: Adding a constraint on the action sampling as $A_{i}\sim P_{\pi,\mathcal{T}}(A^{\texttt{thk}}|E_{i-1},O_{i-1})$ if $A_{i-1}\in\mathcal{A}^{\neg\texttt{thk}}$ else $A_{i}\sim P_{\pi,\mathcal{T}}(A^{\neg\texttt{thk}}|E_{i-1},O_{i-1})$, where $A^{\texttt{thk}}\in\mathcal{A}^{\texttt{thk}}$ and $A^{\neg\texttt{thk}}\in\mathcal{A}^{\neg\texttt{thk}}$ randon varaibles from the partitions of $\mathcal{A}$ standing for the space of textual reasoning and all others, respectively.
  4. 4.
    Reflexion [shinn2023reflexion]: Do multiple trials $\mathcal{F}^{(\leq i)}=\{\mathcal{F}^{(1)},...,\mathcal{F}^{(i)}\}$ until $r^{(i)}:=r(\mathcal{F}^{(i)})=1$ or reach maximum trials, otherwise augmenting the per-trial initial state with an external long-term memory $E^{(i+1)}_{0}=E^{(i)}_{0}\cup\{\mathcal{M}(\mathcal{F}^{(j)},r^{(j)})\}_{j\leq i}$, where the long-term memory $\mathcal{M}(\cdot)$ constructs with textual feedback generated from an LLM prompted by the current trial trajectory and a corresponding binary reward. Thus, the action transition probability, $P_{\pi,\mathcal{T}}(\cdot)$, is continually evolved by reflecting feedback from historical runs.
Overall, this suggests that our agent formulation is general and flexible enough to capture a broad class of representative agentic prompting scenarios.
A.2 Uncertainty Instantiations and Total Uncertainty Expansion.
After defining the agent UQ in Def 2, we presented a trajectory-level total uncertainty that expresses the joint uncertainty across multiple turns in an additive form of the initial query uncertainty, action uncertainty, and observation uncertainty.
Given the joint probability, $P(\mathcal{F}_{\leq T})=P(E_{0},O_{0})\Pi_{i=1}^{T}P_{\pi,\mathcal{T}}(A_{i}|E_{i-1},O_{i-1})P(O_{i}|A_{i},E_{i}),$, we enumerate three example instances of the uncertainty measure $U(\cdot)$ that induces the simple additive form of total uncertainty:
  1. 1.
    Information content (negative log probability), $U(X=x):=-\log P(X=x)$, to measure a point-wise surprisal for a given observation.
  2. 2.
    Entropy, $U(X):=H(X)=\mathbb{E}[-\log P(X)]$, to measure the expected amount of surprisal.
  3. 3.
    Relative entropy, $U(X):=\mathbb{D}_{\rm KL}\big(Q(X)||P(X)\big)$ with a pre-defined reference distribution $Q(X)$ and Kullback-Leibler divergence $\mathbb{D}_{\rm KL}$, to measure the expected amount of surprisal given a prior knowledge (if we have any).
For information content, it is trivial to show by just taking the negative logarithm to $P(\mathcal{F}_{\leq T})$. For the entropy and relative entropy, we have the following chain rule [cover1999elements],
  • $H(\mathcal{F}_{0},...,\mathcal{F}_{T})=\sum_{i=0}^{T}H(\mathcal{F}_{i}|\mathcal{F}_{i-1},...,\mathcal{F}_{1})$,
  • $\mathbb{D}_{\rm KL}\big(Q(\mathcal{F}_{0},...,\mathcal{F}_{T})||P(\mathcal{F}_{0},...,\mathcal{F}_{T})\big)=\sum_{i=0}^{T}\mathbb{D}_{\rm KL}\big(Q(\mathcal{F}_{i}|\mathcal{F}_{i-1},...,\mathcal{F}_{1})||P(\mathcal{F}_{i}|\mathcal{F}_{i-1},...,\mathcal{F}_{1})\big)$,
which directly drives our uncertainty expansion for $U(\mathcal{F}_{\leq T})$ in Section 3.1.
Appendix B Alternative Uncertainty Measures
Although we have mainly focused on the above three information-theoretic uncertainty measures throughout the paper, there are plenty of alternatives one can consider depending on the problem setups. In this section, we examine some possible alternative uncertainty measures, Rényi entropy, Tsallis entropy, and informational energy.
Rényi entropy [renyi1961measures]
has been widely applied in modern machine learning, ecology, biodiversity science, and quantum information, among other fields. Given an order parameter $0<\alpha<\infty$ and $\alpha\neq 1$, it is defined as $H_{\alpha}(X)=\frac{1}{1-\alpha}\log(\sum_{x\in\mathcal{X}}P(x)^{\alpha})$ and generalize many other entropies: max-entropy $H_{0}(X)=\log|\mathcal{X}|$, Shannon entropy $\lim_{\alpha\rightarrow 1}H_{\alpha}(X)=-\sum_{x\in\mathcal{X}}P(x)\log P(x)$, collision entropy $H_{2}(X)=-\log\sum_{x\in\mathcal{X}}P(x)^{2}$, and min-entropy $H_{\infty}(X)\doteq\min_{x\in\mathcal{X}}-\log P(x)$.
Tsallis entropy [tsallis1988possible]
[…]
Onicescu informational energy [pardo1991information, calin2014informational]
is another popular measure of (un)certainty that has bred many applications in quantum mechanics, economics, ecology, social sciences, and so on. It is defined as $\rm{IE}(X):=\sum_{x\in\mathcal{X}}P(x)^{2}$ which can be connected to Rényi entropy of order-2 $H_{2}(X)=-\log\sum_{x\in\mathcal{X}}P(x)^{2}=-\log\rm{IE}(X)$, as well as the power entropy [vajda1968bounds, vajda2007generalized] $V_{2}(X)=1-\sum_{x\in\mathcal{X}}P(x)^{2}=1-\rm{IE(X)}$. There is also a concept, Onicescu’s correlation coefficient [calin2014informational], $\rho(X,Y)=\frac{\sum_{x,y}P(x)P(y)}{\sqrt{\rm{IE(X)\cdot\rm{IE(Y)}}}}$, to quantify the dependency structure between multiple variables.
These sophisticated measures may be worth investigating for the LLM agents’ inference interface, characterized by long-range interactions, evolving memory, and the multifractal property.
Appendix C Extended Context
Connection to the probabilistic Turing machines.
At some level of abstraction, one can view the LLM agents as probabilistic interactive Turing machines (PITMs) [john1977turing] that induce a distribution over action-observation transcripts. This connection may allow us to borrow some possible formal tools from the PITM literature to design UQ method for LLM agents. However, we note the following distinctions: (1) PITMs have an explicit randomness source and a defined distribution, whereas LLM agent also has a procedural uncertainty induced by the decoding time strategies; (2) while PITMs interact via fixed channels, LLM agents commonly interact with users and tools in partially observable or changing formats; and (3) the goal of LLM agents are often underspecified upfront and can be negotiated over turns depending on the situation, though PITMs have a concrete accept/output condition.
Connection to the partially observable Markov decision process.
One might draw an analogy between the agent UQ and belief tracking in Partially Observable Markov Decision Processes (POMDPs) [kaelbling1998planning]. However, the agent theme challenges the standard setups in traditional POMDPs: (1) there is no explicit belief state, but the agent carries an implicit belief in its memory; (2) actions and observations comprise language and structured strings, spanning an effectively infinite space—far from the small discrete spaces in classical POMDP; and (3) it concerns more on the agent’s uncertainty, whereas POMDP concerns mostly on environment uncertainty.
These connections highlight the distinctive edge of agent UQ, while grounding it in a well-established classical problem set.
Appendix D Note on Implementation
While we set our goal to provide a principled vision with a high-level roadmap for agent UQ, we also discuss some implementation recipes in this section to provide useful guidance for practitioners and researchers.
Action classifier.
To realize the conditional information gate, we first need to classify actions to determine whether each action will contribute to total uncertainty reduction for that task or not. As emphasized in Section 3, considering the interactivity of actions is key to accounting for agent uncertainty dynamics, as well as the evidentiality (or factuality) of the action that is solely considered for the existing approaches so far. Therefore, we may want to implement a compound classifier with both interactivity assessment and evidentiality verification. They can be implemented through a rule/syntactic-based, simple verifier [mu2024rule], an LLM judge with specialized prompting, fine-tuning [liu2025compassverifier], or a hybrid of rule-based and LLM-based [peng2025verif].
Action Category Interactivity Scenario-specific Example
Information-gathering interactive Search flight information ($\tau$-bench Airline); retrieve order status ($\tau$-bench Retail); read message box (ToolSandbox)
Asking clarification or confirmation to user interactive Ask for flight choice or request final booking decision ($\tau$-bench Airline); instruct or ask the user to choose cancellation ($\tau$-bench Retail); ask for providing contact information or request updating contact number (ToolSandbox)
Thinking non-interactive Plan flight schedule ($\tau$-bench Airline); consider returning items ($\tau$-bench Retail); think the reason of failed temperature checking (ToolSandbox)
State-changing tool call (write) non-interactive Update flight booking ($\tau$-bench Airline); return/cancel items ($\tau$-bench Retail); Turn on Wifi (ToolSandbox)
Providing final information to user non-interactive Summarize updated flights ($\tau$-bench Airline); summarize returned items ($\tau$-bench Retail); summarize phonebook update results (ToolSandbox)
Table 2: Example action classes in three scenarios from two benchmarks, $\tau$-bench [yao2024tau, barres2025tau] and ToolSandbox [lu2025toolsandbox].
Uncertainty estimation.
In our proposed total uncertainty lower bound Eq. 6, there are three uncertainty terms, including initial query uncertainty, action uncertainty, and observation uncertainty. The action uncertainty can be estimated or approximated by leveraging existing techniques, such as sampling-based ones entropy [malinin2021uncertainty, kuhn2023semantic], pure probability-based methods [marina2020unsupervised, duan2024shifting], hybrid [vashurin2025uncertainty], or the conformal prediction set size [kumar2023conformal].
However, the observation uncertainty and the initial query uncertainty will be much harder to estimate, as we usually do not have full knowledge of the world. If the system allows multiple trials, one can utilize a sampling-consistency-based approach to estimate the observation uncertainty, likewise the traditional black-box UQ methods. If the system does not allow this, one may try to construct a datastore that contains some relevant data from a similar task and leverage nonparameter estimation methods [ren2023outofdistribution, kotelevskii2022nonparametric]. Another line of methods can be establishing a world model to approximate the environment transition and estimate observation uncertainty from it [chae2025web, zhou2025dinowm, anonymous2026dreamphase].
Mutual information estimation.
The proposed information gating mechanism features the information reduction based on the amount of information gain by measuring the conditional mutual information between the current observation $O_{i}$ and initial query $O_{0}$, given all previous conversations except the initial query $E_{i}\backslash O_{0}$. While the estimation of mutual information is notorious for its difficulty, recent advancements on neural estimators [belghazi2018mutual, cheng2020club, mukherjee2020ccmi, molavipour2020conditional, gritsai2025mist] shed light on this problem by enabling estimation in high-dimensional, unstructured data space.
Conditional certainty maximization process.
As an analogy to the conditional uncertainty reduction process presented in Section 4.2, we can also envision a “certainty maximization” approach by replacing the uncertainty with a certainty measure, such as informational energy [pardo1991information], mentioned in Appendix B. Given that Onicescu’s informational energy brings its correlation coefficient as well, we may be able to implement the exact same style information-gating-based method that has the opposite direction to uncertainty reduction dynamics.
Appendix E Action Categorization
Along with the flight booking scenario in Table 1, we extend our action category analysis to two more scenarios, retail and telecommunication tasks, from $\tau$-bench [yao2024tau, barres2025tau] and ToolSandbox [lu2025toolsandbox] in Table 2.
Appendix F Missing Proof
Lemma 2 (Restatement of Lemma 1).
Let the lower bound of agent total uncertainty in Eq. 6 be $\tilde{U}(\mathcal{F}_{\leq T})$, denote $U(X):=H(X)=\mathbb{E}[-\log P(X)]$ and $\text{Info}(X;Y):=I(X;Y)=\mathbb{E}[\log\frac{P(X,Y)}{P(X)P(Y)}]$, then, we have:
$\tilde{U}(\mathcal{F}_{\leq T})$ $\geq H(E_{0},O_{0})-\sum_{i=1}^{T-1}I(O_{i},O_{0}|E_{i}\backslash O_{0})+H(A_{T}|E_{T-1},O_{T-1}),$
$\tilde{U}(\mathcal{F}_{\leq T})$ $\leq H(E_{0},O_{0})+\sum_{i=1}^{T}H(A_{i},O_{i}|E_{i-1},O_{i-1}).$
Proof.
For $i$, if $A_{i}\in\mathcal{A}^{-}$ (resp. $A_{i}\notin\mathcal{A}^{-}$), turn-level uncertainty $U(\mathcal{F}_{i})$ becomes $I(O_{i},O_{0}|E_{i}\backslash O_{0})$ (resp. $U(\mathcal{F}_{i})=U(A_{i}|E_{i-1},O_{i-1})+U(O_{i}|A_{i},E_{i})=H(A_{i},O_{i}|E_{i-1},O_{i-1})$). Thus, if all the intermediate actions are interactive and evidential, e.g., $A_{i}\in\mathcal{A}^{-}$ for all $1\leq i\leq T-1$ (resp. $A_{i}\notin\mathcal{A}^{-}$ for all $1\leq i\leq T-1$), our total uncertainty lower bound $\tilde{U}(\mathcal{F}_{\leq T})$ becomes a monotonic uncertainty reduction process (resp. monotonic propagation process), deriving the above two inequalities. Here, the final termination-turn action should not be interactive, and the observation in this turn is deterministic in our considered setups (Fig. 2), resulting in the accumulation of last action uncertainty $H(A_{T}|E_{T-1},O_{T-1})$. ∎
Appendix G Discussion and Future Work
Does solely establishing the agent UQ framework actually help us to reliably infer on the agent’s performance on a task? Modern LLMs (especially after post-training) are not well-calibrated [tian2023just], so the uncertainty estimates from these ill-calibrated probabilistic models can not be directly used as a reliable performance indicator. Although we present a high-level roadmap for quantifying uncertainty for LLM agent systems, future work should also explore the calibration of LLM agents as well. The connection between calibration and accuracy [park2020calibrated, oh2024towards] of a predictor shows an exciting future work direction for joint optimization of the agent’s problem-solving capability and calibration simultaneously.
Another promising direction for future investigations would be to extend the agent UQ framework to multimodal setups such as graphical user interface agents [nguyen-etal-2025-gui], where extra challenges occur, such as many-to-many correspondence between modalities [chun2025multiplicity]. Defining and establishing a reliable UQ framework for multimodal agents would offer helpful foundations to develop robustness solutions [oh2025understanding, oh2025visual], safety [zhou2025multimodal, chen2025obvious] and more [shi2025towards].